OpenStudy (anonymous):

Simplify f(x) = the quantity 3 x^-2 - 4x plus 1 all over 3x -1. Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(1/3) is undefined. Make sure to show your work.

OpenStudy (anonymous):

do you mean this? \[\LARGE f(x)=\frac{3x^{-2}-4x+1}{3x-1}\]

OpenStudy (anonymous):

yea except the exponent isnt negative

OpenStudy (anonymous):

you wrote it negative ... check it out ! O_O ! ok.. That's nice . ;) hold on a second I'll think ;)

OpenStudy (anonymous):

\[\LARGE f(1)=\frac{3\cdot 1^2-4\cdot 1+1}{3\cdot 1-1}\] \[\LARGE f(1)=\frac{3-4+1}{3-1}\] \[\LARGE f(1)=\frac{0}{2}\] \[\LARGE f(1)=0\] just substitute given value to the X and you'll get the results .... just follow the steps I've shown ;) Good Luck !

OpenStudy (anonymous):

yea i just tottally blanked out to what you just did. can u explain it cus im confused.

OpenStudy (anonymous):

@tmest1992 if you have: f(x)= something...(anything) f(1)= substitute 1 instead of X... example: f(x)=3x+2 ,f(1)=? f(1)=3*1-2 do you understand it ?